Oa November 16 2018 the following call option on FB stock is available: Bid: 139.5S Avk: 139 35 Sire: 3s1 Vol: 37250560 DECEMBER 2018 (EXPIRATION: 12/14) etion Net AskVol at 610 1.75 .00 8. 6.20 s27 And the following put option are available: Sarike Net 430 +1.15420 40/40 86 If FB stock trades at Bid: 139.55 Ask: 139 6 Size 3xl Vol: 37250560 Explain if you will take a position on the call Dec 2018 or the Put stock price to trade around $150 on December 15h 2018 Dec 2018 if you expect the 1) What is time value and intrinsie value of the call option today? 2) You bought the call option. Calculate the profit or loss If FB stock at the end of December is a priced at S106 b. riced at $150
Answer #1

1.
Buy call

Time value=6.20-1.56=4.64
Intrinsic Value=MAX(139.56-138,0)=1.56

2.
Loss per share=6.20
Loss per contract=620

3.
Profit per share=12-6.2=5.8
Loss per contract=580

4.
=138+6.2=144.2